OpenStudy (anonymous):

solve the equation: I 2x+6 I =12 sn* thats not a 12 or 61 there are lines in front

OpenStudy (anonymous):

Solve 2 equations... 2x + 6 = 12 and 2x + 6 = -12

OpenStudy (anonymous):

so x = 3 or x = -9

OpenStudy (anonymous):

and that is the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whenever it's absolute value, you have to solve the equation twice.

OpenStudy (anonymous):

can i ask you another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I 9X-2 I=7

OpenStudy (anonymous):

Set up 2 equations.... 9x - 2 = 7 and 9x - 2 = -7

OpenStudy (anonymous):

an you know there are lines in fornt of the 9 and after the 2 right

OpenStudy (anonymous):

so its 7 , -7

OpenStudy (anonymous):

No, you have to solve each equation

OpenStudy (anonymous):

Add the 2, and then divide by 9 for each equation

OpenStudy (anonymous):

Do you know what I mean?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

You have to solve each equation for x. In order to do that, you first add 2 to both sides.

OpenStudy (anonymous):

On the first equation, if you add 2, you would get 9x = 9. Then divide both sides by 9 and that would give you x = 1

OpenStudy (anonymous):

If you do that for the other equation, first add 2 to both sides. That would give you 9x = -5. Then divide both sides by 9. x = -5/9

OpenStudy (anonymous):

okay i guess i get it alittle i have to practice more i guess

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

thank you so much :)